A stone weigh 200 pounds is suspended from a rope the stone is pulled aside comes to rest when a horizontal force of 130 pounds is applied. What is the tension in the rope? what angle does the rope makes with the horizontal?
1
Expert's answer
2019-05-08T09:36:27-0400




By Newton's law, the body will be at rest if the sum of the projections of forces is equal. From here we get the angle of deviation from the vertical.


"F_T*sin\alpha-F*cos\alpha=0""tg\alpha=F/F_T=130/200=0.65""\alpha=33°"



but we need a tilt angle to the horizontal:


"\beta=90°-\alpha""\beta=57°"

The tension force of the thread is equal to the sum of the projections


"T=F_t*cos\alpha+F*sin\alpha""T=239 pounds"

Answer: 57°, 239 pounds